OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 8, -16, 32, -64, ...

OpenStudy (anonymous):

a_1rn

OpenStudy (anonymous):

can any of you guys help me with my assignment?

OpenStudy (anonymous):

these are the answers available??? an=-1(3)n − 1 an=1(-3)n − 1 an=3(-1)n − 1 an=3(-3)n − 1

OpenStudy (anonymous):

not sure... i thoght the ratio was -2

OpenStudy (anonymous):

personally i thought the geometric sequence was \[a_1r^n=8(-2)^n\]

OpenStudy (anonymous):

A geometric sequence or a geometric progression is represented by \[a _{n}= ar ^{n-1}\] Where n is the number of the term, a is the first term and r is the common ratio. The first term of a geometric sequence is always a, so in your case, that would be 8 Now you need to find 'r' First term = a = 8 Second term = ar = 8r =-16 Third term = ar^2 = 8r^2 = 32 Divide the second term by the first term and you get r = -2 So the correct answer is =8(−2)^n-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that is not an answer -.- though

OpenStudy (anonymous):

\[solve for x: x-6\div5-4x+4\div5=3\]